OpenStudy (anonymous):

Help!! Let f(x)=7x-13. Find f^-1(x). Let f(x)=x^2-16. Find f^-1(x).

OpenStudy (saifoo.khan):

Do you know how to find an inverse?

OpenStudy (anonymous):

Okay thanks..

hartnn (hartnn):

ask if you still need help.

OpenStudy (anonymous):

I'm still confused..what am I finding the inverse to in my problem?? @hartnn

hartnn (hartnn):

ok, let y=7x-13 now to find the inverse, isolate x =... ? then just replace 'x' and 'y' (you can replace before isolating also.)

OpenStudy (anonymous):

-7x=-13-y?? why would there be a y in there??

OpenStudy (anonymous):

and then divide the whole thing by -7, but they're going to be fractions.

hartnn (hartnn):

y= 7x-13 add 13 to both sides, then divide by 7 [Let there be fractions, its ok]

OpenStudy (anonymous):

So basically you set the problem equal to 7x-13, and either make it equal to y or x?? There is no y in the problem? that's why i am confused

OpenStudy (anonymous):

x=-13-y/-7, and then do you want me to solve for x on y=7x-13 or that's what y equals?? So you use x to solve for y and y to solve for x??

OpenStudy (anonymous):

y = f(x).. They are the same thing

OpenStudy (anonymous):

x and y are the same thing??????

hartnn (hartnn):

i considered f(x) = y for simplicity. i'll write out steps so that you can understand. y= 7x-13 y+13 = 7x x = (y+13)/7 now to find inverse function just replace x and y y= (x+13)/7 so your inverse function is \[f^{-1}(x)=\dfrac{x+13}{7}\] got this ?

OpenStudy (anonymous):

f(x) and y are the same thing, so you re-write it as y = 7x-13

OpenStudy (anonymous):

got that part?

OpenStudy (anonymous):

Yes THANK YOU!!!!

OpenStudy (anonymous):

swop d x for d y.... it becomes x = 7y-13

hartnn (hartnn):

can you do same thing for f(x)=x^2-16. ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

after swopping, solve for y

OpenStudy (anonymous):

7y= x+13 y = (x+13)/7

OpenStudy (anonymous):

Yes x=4 + sqrt y

OpenStudy (anonymous):

yes, but u have to swop d x and d y back... final answer would be y = 4 + sqrt(x)

hartnn (hartnn):

did u simplify like this ? \[\sqrt{16+y}=4+\sqrt y \] no, thats incorrect.

OpenStudy (anonymous):

by d do you mean the?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

sqrt. y+16=x^2 (Root over y+16)

hartnn (hartnn):

yeah, the inverse function will be \[f^{-1}(x)=\sqrt{x+16}\] no further simplification possible.

hartnn (hartnn):

ok ? got it ?

OpenStudy (anonymous):

Oh got it!!!